if the ratio of two supplementary angles is 7 to 8, calculate the measure of the larger of the two angles
Answer by nyc_function(2741)

Let 7x = small angle
let 8x = big angle
7x + 8x = 180
15x = 180
x = 180/15
x = 12
Small angle = 7(12) = 84
Big angle = 8(12) = 96
NOTE: 96 degrees is the only supplementary angle here. Angle 84 degrees is not supplementary but together they add up to 180 degrees.
I hope this helps.